Using the information given in the following statement, find the measures of angles A, B, and C for △ABC.
Here's the information:
m∠B = 3m∠A, m∠C = 2 m∠A + m∠B
(by the way....this is COMPLETELY retarded. 2nd yr in college, final Math class and even the instructor said that it's likely that I will never use anything I learn in this class, in real life. Our education system is genius! NOT!)Thanks - Michael -


Answer by scott8148(6628)

learning the math is really secondary, per se
math is the tool to teach you how to reason and logically solve problems

A + B + C = 180º

B = 3A

C = 2A + B

substituting ___ C = 2A + (3A) = 5A

substituting ___ A + (3A) + (5A) = 180º ___ 9A = 180º ___ A = 20º

substituting ___ B = 3(20º) = 60º

substituting ___ C = 5(20º) = 100º